The pressure on a 205mL volume of gas is decreased from 985mm Hg to 615 mm Hg while constant temperature is maintained. What is the new volume of the gas


1
Expert's answer
2020-07-13T07:23:17-0400

Solution:

by the law of Boyle-Marriott:

p1V1=p2V2, V2 = p1V1/p2 = 2.05*10-4m3*985/615 = 3.2833*10-4m3

Answer: 3.2833*10-4m3